# Volumetrically Consistent 3D Gaussian Rasterization

### v2

@@ 6.2 Tomography
where and are the source and detected intensities.
Using 3DGS out-of-the-box for tomography reconstruction produces poor results. This is because the splatting process isn’t physically-accurate, as shown by -Gaussian Zha et al. [55]. They identify that unnormalized 3D Gaussians (as used by 3DGS), when splatted onto the screen, are view inconsistent. To fix this, they devise a specialized rendering model for tomography by normalizing the 3D Gaussians, which achieves state-of-the-art reconstruction results. Tomography reconstruction is slightly simpler than view synthesis since the forward rendering model does not depend on the ordering of the Gaussians, and there is no emission or color involved, only the transmittance along a ray is measured. While Zha et al. [55] remove the integration bias from the alpha computation process by normalizing Gaussians, their approach still incurs the affine approximation in 3DGS.
Our method requires no changes to the alpha computation process and still works out of the box for tomography. We implement our method for tomography by integrating our rasterizer in Zha et al.’s codebase and evaluate on their real and synthetic datasets. Since our method provides a more accurate forward rendering model for tomography, we produce fewer points compared to Zha et al. during optimization, while still matching them in performance on 3D PSNR and 3D SSIM on both real and synthetic datasets. We show results on their most challenging sparse-view (25 views) reconstruction in Tab. 3, and visual reconstructions from our method and Zha et al. [55] in Appendix F.
@@ 7 Limitations
since our method, like 3DGS, is rasterization-based, it inherits some of its issues. Per-tile (instead of per-pixel) primitive sorting and no overlap handling introduces popping artifacts. Several recent works [39, 21, 33] ameliorate this issue for 3DGS (their methods readily apply to ours), but they cannot fully eliminate it like ray-tracing-based methods can [31].
@@
in cases of large errors in camera calibration or distortion, our method produces opaque artifacts while 3DGS degrades more gracefully by producing blurry reconstructions, as shown in Fig. 6.
@@ 9 Conclusion
We introduce a method that is a more physically-accurate version of 3DGS by computing alpha values via analytical 3D Gaussian integration instead of splatting. Our approach integrates the strengths of physically-based ray tracing with the efficiency of rasterization. It enables higher-quality reconstruction than 3DGS (as measured by SSIM and LPIPS). We also demonstrate that since our method is a plug-and-play replacement for alpha computation in 3DGS, it can be directly used for other 3DGS-based applications such as tomography.
10 Acknowledgements
This work was supported in part by NSF grants 2341952 and 2105806. We also acknowledge ONR grant N00014-23-1-2526, gifts from Adobe, Google, Qualcomm and Rembrand, support from the Ronald L. Graham Chair, and the UC San Diego Center for Visual Computing. We also thank Alexander Mai, Kaiwen Jiang, and Zachary Novack for helpful discussions and perspectives.
@@ 10 Acknowledgements
The appendix is organized as follows. In section Appendix A, we show a qualitative comparison between our volumetric rasterizer and ray marching to visualize the impact of the sorting and non-overlapping assumptions we make in our method (see Sec. 4.1 in the main paper). We derive an expression for the 1D Gaussian distribution along the camera ray in Appendix B (eq.15 in the main paper). We describe the hyperparameters we use and more details about our modifications to 3DGS’s densification strategy in Appendix C (Sec. 5 of the main paper). In Appendix D, we discuss potential reasons for the slowdown in training time for our method (Sec 6.1 of main paper). We report the reconstruction quality metrics for all the scenes individually in Appendix E Finally, in Appendix F, we present visual results for tomographic reconstruction from our method and the current splatting-based state-of-the-art (SOTA) Zha et al. [55], corresponding to the quantitative metrics reported in Sec. 6.2 (Tab. 2) of the main paper.
@@ Appendix A Comparing volumetric rasterizer with ray marching
Our method enables volumetrically consistent rasterization by analytically computing the volume rendering integral for each primitive. Our approach is accurate under the assumption of correctly sorted and non-overlapping primitives (Sec. 3.2 main paper). In this section, we qualitatively show the effect of this approximation, i.e. to what extent and in which situations it deviates from a reference solution to the volume rendering equation without approximations.
Consequences of overlapping primitives have no relation to view dependence of color, so we assume view-independent colors in this experiment to simplify the ray-marching process.
@@
The 1D distribution along the ray is given by Eq. 15:
@@
Appendix C Implementation details and hyperparameters
@@
Unlike 3DGS which uses 2D screen space positional gradients, we use the world space 3D positional gradients to compute the gradient norm for densification thresholding, similar to [36, 31]. We use separate positional gradient thresholds for the splitting and cloning operations. We require much lower gradient threshold values since our method produces more opaque primitives than 3DGS which reduces the positional gradient value. We densify and prune every 200 iterations. For pruning, we prune based on the value of , the parameter used to parameterize the density of the primitives (see eq. 20 in the main paper). We remove primitives for which is less than a minimum opacity threshold. We will open-source the code after publication.
@@
The learning rates for density and color features are set to and respectively. Initial and final position learning rates are set to and . For our method, we adjust the parameter defined as percent_dense in the 3DGS code base. We set it to , and for the Mip-NeRF360, Tanks&Temples, and DeepBlending datasets respectively. We adjust the (hardcoded) pruning hyperparameter to prune points with large scales to 0.01 from the default value 0.1.
Appendix D Training times
We use a Nvidia 3090 Ti for training both our method and 3DGS. Our training times are similar to 3DGS, with a slight slowdown that arises from the extra computations needed for our alpha computation compared to 3DGS. 3DGS can re-use the splatted 2D Covariance matrix to compute alpha for each pixel. Our method requires computing and other intermediaries for each pixel separately, leading to a slight slowdown. We compute inverses of both the 2D and 3D covariances, as opposed to 3DGS which does the inverse computation only once for the 2D covariance in the vertex shader phase. The vertex shader in our slang.D implementation uses atomic add operations to write the computed inverse covariance matrices to global memory. Our method requires more atomic adds per primitive to store 3D covariance matrices, leading to slower performance compared to 3DGS, which requires fewer atomic adds to store 2D covariance matrices.
Appendix E Reconstruction metrics per scene
In the main paper (Tab. 1), we report metrics averaged over all scenes for the three datasets -Mip-NeRF360, Tanks&Temples, and DeepBlending. We report metrics from our method for each scene individually in Mip-NeRF360, Tanks&Temples, and DeepBlending datasets in Tab. 4, Tab. 5 and Tab. 6. For per-scene metrics on some of the other baselines, please see [25, 19].
@@ Appendix E Reconstruction metrics per scene
Appendix F Tomography reconstruction
We present tomography reconstructions for our method and -Gaussian (Zha et al. [55]), corresponding to quantitative results referenced in Tab. 2, Sec. 6.2 of the main paper. We demonstrate comparable performance to [55], which is the current state-of-the-art splatting-based tomography method. We observe similar visual quality for both methods (see Fig. 8), with our method producing fewer primitives (Sec. 6.2 in the main paper). Recall that our approach works out of the box, modeling the forward rendering tomography process more accurately than Zha et al. [55]. While [55] modifies the rendering algorithm to account for integration bias, (Sec. 6.2 in the main paper), it still retains the affine approximation in splatting, referenced in Sec. 3.2 of the main paper.

### v3

@@ 6.2 Tomography
where is the source intensity and is the detected intensity.
Using 3DGS out-of-the-box for tomography reconstruction produces poor results. This is because the splatting process isn’t physically accurate, as shown by -Gaussian Zha et al. [58]. They identify that unnormalized 3D Gaussians (as used by 3DGS), when splatted onto the screen, are view inconsistent. To fix this, they devise a specialized rendering model for tomography by normalizing the 3D Gaussians, which achieves state-of-the-art reconstruction results. Tomography reconstruction is slightly simpler than view synthesis since the forward rendering model does not depend on the ordering of the Gaussians, and there is no emission or color involved, only the transmittance along a ray is measured. While Zha et al. [58] remove the integration bias from the alpha computation process by normalizing Gaussians, their approach still incurs the affine approximation in 3DGS.
Our method requires no changes to the alpha computation process and still works out of the box for tomography. We implement our method for tomography by integrating our rasterizer in Zha et al.’s codebase and evaluate on their real and synthetic datasets. Since our method provides a more accurate forward rendering model for tomography, we produce fewer points compared to Zha et al., while still matching them in performance on 3D PSNR and 3D SSIM on both real and synthetic datasets. We show results on their most challenging sparse-view (25 views) reconstruction in Tab. 2 and a subset of visual results in Fig. 5.
@@ 7 Limitations
Since our method, like 3DGS, is rasterization-based, it inherits some of its issues. Per-tile (instead of per-pixel) primitive sorting and no overlap handling introduces popping artifacts. Several recent works [42, 23, 36] ameliorate this issue for 3DGS (their methods are readily applicable to us), but they cannot fully eliminate it like ray-tracing-based methods can [34].
@@
In cases of large errors in camera calibration or distortion, our method produces opaque artifacts while 3DGS degrades more gracefully by producing blurry reconstructions, as shown in Fig. 6. Refer to supp.I for a thorough analysis of the failure cases in the MipNerf-360 dataset, which exemplifies this effect.
@@ 9 Conclusion
We introduce a more physically-accurate version of 3DGS by computing alpha values via analytical 3D Gaussian integration instead of splatting. Our approach integrates the strengths of physically-based ray tracing with the efficiency of rasterization. It enables higher quality reconstruction than 3DGS (as measured by SSIM and LPIPS). We also show that since our method is a plug-and-play replacement for alpha computation in 3DGS, it can be directly used for other 3DGS-based applications such as tomography.
Acknowledgements:
This work was supported in part by NSF grants 2341952 and 2105806. We also acknowledge ONR grant N0001423-1-2526, gifts from Adobe, Google, Qualcomm and Rembrand, support from the Ronald L. Graham Chair, and the UC San Diego Center for Visual Computing. We also thank Alexander Mai, Kaiwen Jiang, and Zachary Novack for helpful discussions and perspectives.
@@ Acknowledgements:
The appendix material is organized as follows. In section Appendix A, we show a qualitative comparison between our volumetric rasterizer and ray marching to visualize the impact of the sorting and non-overlapping assumptions we make in our method (see Sec. 4.1 in the main paper). We derive an expression for the 1D Gaussian distribution along the camera ray in Appendix B (eq.15 in the main paper). In Appendix C we validate that our method outperforms 3DGS at approximating sharp edges and constant texture regions in a multiview setting on the nerf-synthetic dataset (Sec. 4 in the main text). We describe the hyperparameters we use and more details about our modifications to 3DGS’s densification strategy in Appendix D (Sec. 5 of the main paper), and report the reconstruction quality metrics for all the scenes individually in Appendix E. We report our average training times in Appendix F (Sec 6.1 of main paper). In Appendix G we compare the reconstruction quality (SSIM and LPIPS) of 3DGS and our method across different memory budgets, referenced in Sec. 6.1 of the main paper. We present the results of an ablation study on our method in Appendix H. We analyze the failure cases from our method compared to 3DGS on the MipNeRF-360 dataset (see Sec. 7 in the main text) and present a quantitative analysis in Appendix I.
@@ Appendix A Comparing volumetric rasterizer with ray marching
Our method enables volumetrically consistent rasterization by analytically computing the volume rendering integral for each primitive. Our approach is accurate under the assumption of correctly sorted and non-overlapping primitives (Sec. 3.2 main paper). In this section, we qualitatively show the effect of this approximation, i.e. to what extent and in which situations it deviates from a reference solution to the volume rendering equation without approximations. Consequences of overlapping primitives have no relation to view dependence of color, so we assume view-independent colors in this experiment to simplify the ray-marching process.
@@
The 1D distribution along the ray is given by (eq. 15 main paper):
@@
Appendix C Comparing our method and 3DGS on nerf-synthetic
Recall that in Fig.2 and Fig.3 in the main text, we fit a single image with a fixed number of primitives and observe that our method is better than 3DGS at approximating sharp edges and constant regions. Here, we validate this observation in a multi-view setting on the nerf-synthetic (NS) dataset. We optimize our method and 3DGS with the same initialization, and no densification to isolate the expressiveness of our method from the optimization dynamics. For all metrics (PSNR, SSIM, LPIPS) we (29.81, 0.941, 0.074) outperform 3DGS (29.56, 0.936, 0.082) averaged over all 8 scenes in NS. Our method renders piece-wise constant textures (common in nerf-synthetic) better than 3DGS. See Fig. 8 for a qualitative comparison.
Appendix D Implementation details and hyperparameters
@@
Unlike 3DGS which uses 2D screen space positional gradients, we use the world space 3D positional gradients to compute the gradient norm for densification thresholding, similar to [39, 34]. We use separate positional gradient thresholds for the splitting and cloning operations. We require much lower gradient threshold values since our method produces more opaque primitives than 3DGS which reduces the positional gradient value. We densify and prune every 200 iterations. For pruning, we prune based on the value of , the parameter used to parameterize the density of the primitives (see eq. 20 in the main paper). We remove primitives for which is less than a minimum opacity threshold. Similar to [34], we apply softplus activation () to the spherical harmonic coefficients.
@@
The learning rates for density and color features are set to and respectively. Initial and final position learning rates are set to and . In the 3DGS codebase, the scene extent (or camera extent) is defined as the maximum distance between any training camera and the average camera centroid. The scene extent controls the splitting, cloning, and pruning thresholds in the densification process. In our method, we scale the scene extent by a factor of , and for the Mip-NeRF360, Tanks&Temples, and DeepBlending datasets respectively. We also adjust the (hardcoded) pruning hyperparameter to prune points with large scales to 0.01 from the default value 0.1.
Appendix E Per scene metrics
@@
We report metrics from our method for each scene individually in Mip-NeRF360, Tanks&Temples, and DeepBlending datasets in Tab. 3, Tab. 4 and Tab. 5. Note that in the main paper (Tab. 1), we report metrics averaged over all scenes for the three datasets -Mip-NeRF360, Tanks&Temples, and DeepBlending. For per-scene metrics on some of the other baselines, please see [28, 21].
Appendix F Average training times
We use a Nvidia 3090 Ti for training both our method and 3DGS. Our training times are similar to 3DGS, with a slight slowdown that arises from the extra computations needed for our alpha computation compared to 3DGS. 3DGS is able to re-use the splatted 2D Covariance matrix for computing alpha for each pixel. Our method requires computing and other intermediaries for each pixel separately which leads to a slight slowdown. We compute inverses of both the 2D and 3D covariances, as opposed to 3DGS which does the inverse computation only once for the 2D covariance in the vertex shader phase. The vertex shader in our slang.D implementation uses atomic add operations to write the computed inverse covariance matrices to global memory. Our method requires more atomic adds per primitive to store 3D covariance matrices, leading to slower performance compared to 3DGS, which requires fewer atomic adds to store 2D covariance matrices.
For MipNeRF-360 scenes, on average 3DGS and our method generate 2.62M and 2.93M points respectively, and the training times for 30000 iterations are 51.5 min and 61 min respectively. Similarly for Tanks&Temples, on average 3DGS and our method generate 1.79M and 1.72M points respectively, and the training times for 30000 iterations are 20.5 min and 33.2 min respectively.
ZipNerf training:
We trained and evaluated ZipNeRF [6] on a Nvidia A-40 GPU, using their official code release [7]. In the official code release for ZipNeRF, the indoor and outdoor scenes in the Mip-NeRF360 dataset are evaluated at 2x and 4x downsampled resolutions respectively. We trained and evaluated ZipNeRF at full resolution on the Tanks&Temples and DeepBlending datasets, but for the MipNeRF-360 dataset, we used 2x downsampled images (for both indoor and outdoor scenes) to fit them in the GPU memory. ZipNeRF took approximately 12-13 hours per scene to train on our GPU. For the Tanks&Temples dataset, we re-ran COLMAP before training Zip-NeRF as highlighted in issue 7 in the official Zip-NeRF code release [7].
Appendix G Reconstruction quality for different memory budgets
In Fig. 9 we compare the performance of our method against 3DGS across different memory budgets, i.e. across a different number of maximum primitives that each method is allowed to generate. We stop densification and pruning once the optimization reaches the maximum number of allowed primitives. For each memory budget, we average the results over 4 scenes, kitchen, stump, train, counter. For the same memory budget, our method consistently outperforms 3DGS on both SSIM and LPIPS. This holds across a range of memory budgets as shown in Fig. 9. Recall that our method represents opaque textures better than 3DGS for the same number of primitives. We demonstrated this through the toy examples in Fig. 2 and Fig. 3 in the main paper. Since LPIPS and SSIM measure perceptual similarity and edge quality, this also translates to quantitative improvements as demonstrated in Fig. 9.
Appendix H Ablation study
| PSNR | SSIM | LPIPS | Points | |
| 3DGS | 21.38 | 0.588 | 0.360 | 3.4M |
| 3DGS + Our HyperParams | 20.66 | 0.571 | 0.329 | 7.4M |
| Ours w/o Reparam | 21.23 | 0.600 | 0.309 | 7M |
| Ours w EVER Reparam | 20.04 | 0.519 | 0.396 | 9.4M |
| Ours | 21.41 | 0.615 | 0.306 | 3.8M |
We conduct an ablation study to assess the impact of hyperparameters and density parameterization on our method, with results shown in Tab. 6. We compare our method with 3DGS (row 1), and for fairness, we also run 3DGS with the hyperparameters used by our method. Using 3DGS with our hyper-parameters (row 2) improves LPIPS, but worsens PSNR and SSIM while generating almost 2x more primitives compared to the base 3DGS configuration (row 1). Our method (row 4) fares relatively better on all metrics.
We also run our method without density reparameterization (row 3); we ensure positive density by applying the softplus activation. We observe improvement in LPIPS and SSIM compared to 3DGS but at almost double the number of primitives. We also experiment with the parameterization proposed in Mai et al. [34] (row 4), which results in poorer SSIM and LPIPS while producing too many points compared to our method (row 5).
| Dataset | Worst Gap LPIPS | Better / Comparable % | Failure % | Total Images |
| flowers | -0.022 | 100.0 | 0.000 | 21 |
| room | 0.005 | 97.4 | 2.632 | 38 |
| bonsai | 0.004 | 100.0 | 0.000 | 36 |
| bicycle | -0.008 | 100.0 | 0.000 | 24 |
| treehill | 0.017 | 88.2 | 11.765 | 17 |
| counter | -0.006 | 100.0 | 0.000 | 29 |
| stump | 0.004 | 100.0 | 0.000 | 15 |
| garden | 0.007 | 91.3 | 8.696 | 23 |
| kitchen | 0.004 | 100.0 | 0.000 | 34 |
| Dataset | Worst Gap SSIM | Better / Comparable % | Failure % | Total Images |
| flowers | 0.005 | 100.0 | 0.000 | 21 |
| room | -0.004 | 100.0 | 0.000 | 38 |
| bonsai | -0.006 | 100.0 | 0.000 | 36 |
| bicycle | -0.002 | 100.0 | 0.000 | 24 |
| treehill | -0.035 | 58.8 | 41.176 | 17 |
| counter | -0.001 | 100.0 | 0.000 | 29 |
| stump | -0.013 | 93.3 | 6.667 | 15 |
| garden | -0.005 | 100.0 | 0.000 | 23 |
| kitchen | -0.007 | 100.0 | 0.000 | 34 |
| Dataset | Worst Gap PSNR | Better / Comparable % | Failure % | Total Images |
| flowers | -1.204 | 95.2 | 4.762 | 21 |
| room | -1.152 | 97.4 | 2.632 | 38 |
| bonsai | -2.328 | 94.4 | 5.556 | 36 |
| bicycle | -0.414 | 100.0 | 0.000 | 24 |
| treehill | -1.174 | 88.2 | 11.765 | 17 |
| counter | -2.163 | 93.1 | 6.897 | 29 |
| stump | -0.636 | 100.0 | 0.000 | 15 |
| garden | -0.536 | 100.0 | 0.000 | 23 |
| kitchen | -5.764 | 91.2 | 8.824 | 34 |
Appendix I MipNeRF-360 dataset failure case analysis
We analyze failure cases for the 3 metrics. We define a failure case as being worse than 3DGS by a threshold, separately chosen for each metric (1dB PSNR, 0.01 SSIM, 0.005 LPIPS). Averaged on all test images in the dataset, our failure rate is 4.64%, 3.38%, and 2.11% for PSNR, SSIM and LPIPS respectively. treehill shows the highest failure rate among all scenes, with LPIPS and SSIM differences of 0.017 and 0.027, respectively (example in Fig. 10). For PSNR, we observe an outlier failure case from kitchen, 5.6 dB worse than 3DGS (Fig. 10). However, averaged over all 34 kitchen test images, our method (31.51 dB) performs only slightly worse than 3DGS (31.67 dB) in PSNR. See Tab. 7 for per-scene quantitative analysis on all three metrics for Mip-NeRF360 dataset.